Four sisters are all two years apart in age an the youngest is 4 years old. In how many years will the four daughters together be as old as their father, who is 46 now?
Answer by math_helper(2461)


Their current ages are 4, 6, 8, and 10 ( 2 yrs apart, youngest is 4)

Let x be the number of years until (4+x)+(6+x)+(8+x)+(10+x) = 46+x
28 + 4x = 46+x
3x = 18
x = 6

Ans: In 6 years the four sister's ages will add up to their father's age.

Check:
Sisters: (4+6) + (6+6) + (8+6) + (10+6) = 10 + 12 + 14 + 16 = 52
Father: 46 + 6 = 52